2x^2-4x-1=0


answer: 2 + sqrt6 /2 , or 2 - sqrt6/2 anyone can give me a full break down of why trhe answer is 2 + sqrt 6/2 or 2 - sqrt6/2

Answer by stanbon(75887)

2x^2-4x-1=0
This is a quadratic with a = 2 ; b = -4 ; and c = -1
---
Use the Quadratic Formula:
x = [-b +- sqrt(b^2 - 4ac)]/(2a)
-------------------------------------
Plug in the values for a, b, and c to get:
x = [4 +- sqrt(16 - 4*2*-1)]/4
---
x = [4 +- sqrt(24)]/4
---
x = [4 +- 2sqrt(6)]/4
---
x = (1 + (1/2)sqrt(6)) or x = (1- (1/2)sqrt(6))
This can be written:
x = (2+sqrt(6))/2 or x = (2-sqrt(6))/2
=====================================================
Cheers,
Stan H.